# A semi-supervised framework for diverse multiple hypothesis testing scenarios

### v2

@@ Abstract
Standard multiple testing procedures are designed to report a list of discoveries, or suspected false null hypotheses, given the hypotheses’ p-values or test scores. Recently there has been a growing interest in enhancing such procedures by combining additional information with the primary p-value or score. Specifically, such so-called side information can be leveraged to improve the separation between true and false nulls along additional dimensions thereby increasing the overall sensitivity. In line with this idea, we develop RESET (REScoring via Estimating and Training), which uses a unique data-splitting protocol that subsequently allows any semi-supervised learning approach to factor in the available side information while maintaining finite-sample error rate control. Our practical implementation, RESET Ensemble, selects from an ensemble of classification algorithms so that it is compatible with a range of multiple testing scenarios without the need for the user to select the appropriate one. We apply RESET to both p-value and competition based multiple testing problems and show that RESET is (1) power-wise competitive, (2) fast compared to most tools and (3) able to uniquely achieve finite sample false discovery rate or false discovery proportion control, depending on the user’s preference.
Corresponding authors: J.Freestone@maths.usyd.edu.au, uri.keich@sydney.edu.au
@@ 1 Introduction
Scientists are often interested in testing many null hypotheses, $\{H_{i}:i=1,\dots,m\}$. As an example, a scientist may be interested in assessing which genes exhibit a change in gene expression levels in response to a particular drug treatment. In this case, the $i$th null hypothesis $H_{i}$ states that ‘the $i$th gene exhibits no change in gene expression level’. This type of large-scale hypothesis testing is usually achieved by assigning a p-value, $p_{i}$, to each null hypothesis. The collection of p-values then undergoes a filtering process that rejects a subset of these null hypotheses that have sufficiently small p-values subject to a type-1 error rate control to minimize the number of misreported true null hypotheses.
@@
The most common choice of error rate in this setup is the false discovery rate (FDR), which is defined as the expectation of the false discovery proportion (FDP) — the fraction of true null hypotheses in the reported list of rejections, or discoveries [2]. In this case, the filtering procedure ensures that the FDR is $\leq\alpha$ where $\alpha\in\left(0,1\right)$ is a prespecified threshold. Alternatively, we might be interested in controlling the FDP rather than its expectation in the following sense. Given a confidence parameter $\gamma$ and a threshold $\alpha$, the FDP is said to be controlled if $\mathbb{P}(FDP>\alpha)\leq\gamma$. Evidently, such FDP control reduces the risk that the realized FDP exceeds $\alpha$, which is not guaranteed by FDR control [40, 27, 19, 38, 16, 20].
Determining a p-value for each null hypothesis can be challenging because (a) it requires knowledge of the underlying null distribution and (b) it can be computationally intensive. A recently developed competition framework offers an alternative approach to FDP/FDR control that does not require computation of canonical p-values [1, 21]. Instead, all that these procedures require is a single draw, $\tilde{Z}_{i}$, from each null distribution. The null drawn $\tilde{Z}_{i}$ is then compared with $Z_{i}$, the originally observed score (test statistic) for $H_{i}$, thus conceptually providing a crude 1-bit p-value.
These p-values are too crude to be used in typical p-value based filters. Instead, assuming that larger values of $Z_{i}$ provide more evidence against $H_{i}$, we compete each pair of scores $(Z_{i},\tilde{Z}_{i})$ by recording a label $L_{i}\in\{\pm 1\}$ indicating which was the largest of the two scores, as well as a winning score, $W_{i}=f(Z_{i},\tilde{Z}_{i})$. The function $f$ needs to be symmetric in its arguments and should increase with the values of $Z_{i}$ and $\tilde{Z}_{i}$, e.g., $W_{i}=Z_{i}\vee\tilde{Z}_{i}$, or with the discrepancy between the two, e.g., $W_{i}=\lvert Z_{i}-\tilde{Z}_{i}\rvert$. [^1] In general, the functions can be hypothesis-dependent, i.e., $f_{i}$ instead of $f$. For each true null hypothesis, $L_{i}$ is equally likely to be ±$\pm 1$ independent of all the scores $W=\{W_{i}\}$ and all other labels $L_{-i}$ because both scores were sampled from the same null distribution and $f$ is symmetric. Hence, the number of negative labels, those with $L_{i}=-1$, can be used to estimate the number of true nulls with positive labels among the top scoring hypotheses. Using this fact, the labels and scores $\{(L_{i},W_{i}):i=1,\dots,m\}$ are passed through a filter to reject a subset of the top scoring hypotheses, with either the FDR or FDP controlled depending on the user’s preference.
@@
More recently, research has been focused on improving these filtering methods by leveraging side information $\mathbf{x}_{i}\in\mathcal{X}$ that complements each p-value $p_{i}$ or score $W_{i}$ to better discriminate between the true and false null hypotheses. These methods utilize a variety of strategies but ultimately operate with the same goal in mind: the side information is used to weight or rescore the hypotheses according to how confidently we believe that each hypothesis is a false null. Indeed, there are several such tools that augment p-values with side information [25, 34, 33, 37, 51, 35]. When it comes to using side information in the competition framework, we are aware of only one fully specified method with rigorous error-rate control, which is the recently published Adaptive Knockoffs [44] (AdaPT [33] only offers a meta-procedure).
@@
In this paper, we offer a novel approach for multiple testing with side information. The idea is to randomly split a set of suspected true nulls into two sets, where one set is used for training a semi-supervised learning model to discriminate between the true and false null hypotheses, and the other for estimating the number of false discoveries. Accordingly, we refer to our new approach as RESET—REScoring via Estimating and Training. Our method flexibly incorporates any semi-supervised approach to help distinguish between true and false null hypotheses. Moreover, it is equally applicable to the competition and p-value context and allows the user to choose between FDR or FDP control. We evaluate RESET in a range of simulations and real data applications to verify that it is a competitive alternative to existing methods in terms of statistical power and speed.
@@ 2 Background: competition-based multiple testing
Competition-based multiple testing was first established in the proteomics community where the goal is to infer which proteins, i.e., long chains of amino acids, are present in a biological sample [17]. Direct protein detection is challenging, and so a bottom-up approach is taken where the proteins are broken into small chains of amino acids, called peptides, and the objective is to identify these peptides instead. Then, a process called liquid chromatography tandem mass spectrometry (LC-MS/MS) is used to isolate each sample peptide and subsequently generate a fragmentation spectrum. Each such spectrum can be thought of as a fingerprint of the peptide that generated the spectrum, allowing us to hypothesize which peptides are present in the sample.
As outlined in the introduction, we accompany $H_{i}$ with a score $Z_{i}$, defined as the maximum scoring PSM associated with the $i$th target peptide, and the higher the score, the more likely $H_{i}$ is false, i.e., the peptide is in the sample. Similarly, we define $\tilde{Z}_{i}$ as the maximum scoring PSM associated to the $i$th target’s corresponding decoy peptide. For each true null hypothesis, we expect $Z_{i}$ and $\tilde{Z}_{i}$ to be drawn from the same (hypothesis-specific) null distribution, independently of everything else. Finally, we obtain a winning score $W_{i}=Z_{i}\vee\tilde{Z}_{i}$ and a label $L_{i}$ indicating whether the winning score was from the $i$th target or its decoy [42].
Subsequently, the statistics community independently formalized this idea in the context of variable selection, where the goal is to identify relevant variables from $\{X_{i}:i=1,\dots,p\}$ that are associated with a response $y$. The procedure that achieves this while controlling the FDR is called the knockoff filter, which begins by pairing each variable $X_{i}$ with an artificial variable $\tilde{X}_{i}$, called a knockoff. In this paper, we consider the model-X version of the knockoff filter [7] which makes the following assumptions [^2] However, our results are equally applicable to the fixed-X model as well [1].. First, the joint distribution of the variables, $X$, is assumed to be known. Second, each observation $(X_{i,1},\dots,X_{i,p},y_{i})$ is sampled i.i.d. from the joint distribution $F_{XY}$. Because $X$ is random and its distribution known, the knockoffs $\tilde{X}$ are also constructed randomly subject to the following conditions:
  1. 1.
[…]
  2. 2.
    We construct $\tilde{X}$ independently of $y$ by only looking at $X$, i.e., $\tilde{X}\perp\!\!\!\perp y\mid X$.
Once the knockoffs have been constructed, each variable is given a user-defined score $W_{i}:=W_{i}([X,\tilde{X}],y)\in\mathbb{R}$ and a label $L_{i}([X,\tilde{X}],y)\in\{\pm 1\}$. The idea is that the label $L_{i}$ indicates whether the variable $X_{i}$ or the knockoff $\tilde{X}_{i}$ has greater evidence for being relevant to the model, and the score $W_{i}$ indicates by how much. Commonly, the score and label are combined by setting $sign(W_{i})=L_{i}$; however, as will become clear in our later presentation of RESET, it is instructive to separate them.
In this paper, we consider the Lasso Coefficient Difference (LCD) score, which is the difference in the magnitude of the coefficients of the Lasso regression [49]. Specifically, cross-validation is used to determine the $\lambda$ penalty when regressing on the augmented design matrix $[X,\tilde{X}]$. We denote the absolute values of the Lasso($\lambda$)-fitted coefficients as $Z_{i}$ for the variables and $\tilde{Z}_{i}$ for the knockoffs. The scores and labels are then defined as
$W_{i}:=\big{|}Z_{i}-\tilde{Z_{i}}\big{|},\quad L_{i}:=sign(Z_{i}-\tilde{Z_{i}}).$ (1)
Hypotheses with zero scores are typically ignored. In either context, peptide inference or variable selection, we expect that the labels of the true null hypotheses are independently and equally likely to be ±$\pm 1$, independently of everything else. This is formalized in the next assumption.
@@
Let $N$ be the indices of the true null hypotheses. The labels $\{L_{i}:i\in N\}$ are i.i.d. ±$\pm 1$ uniform random variables independent of all the scores $W$ and the labels of the false null hypotheses.
For the knockoff filter, Assumption 1 follows from the knockoff construction, whereas in the peptide detection problem, Assumption 1 is based on our previous discussion that the scores of a null target and its decoy, $Z_{i}$ and $\tilde{Z}_{i}$, are independent draws from the same null distribution, and moreover, this happens independently of all other pairs of scores. This assumption is commonly applied in the literature, with several analyses validating this claim empirically [39, 21].
Given the above assumption, the number of hypotheses with $L_{i}=-1$ (i.e., decoy or knockoff wins) that score above a score cutoff of $\tau$ can be used to estimate the number of true null hypotheses that score above $\tau$ with $L_{i}=1$. Using this strategy, we can report a list of discoveries while satisfying a type-1 error rate control in one of two ways. The first and by far the more common way is to apply Selective SeqStep+ (SSS+) [1], which controls the FDR at a prespecified threshold $\alpha$ by reporting the following list of discoveries $R_{\alpha}$:
$R_{\alpha}:=\{i\in[p]:W_{i}\geq\tau,L_{i}=1\},\quad\text{where }\tau:=\min\left\{t\in\mathbb{R}:\frac{\#\{W_{i}\geq t,L_{i}=-1\}+1}{\#\{W_{i}\geq t,L_{i}=1\}\vee 1}\leq\alpha\right\}.$ (2)
SSS+ is made precise in Algorithm S1 and works in a more general setting when the probabilities of the positive and negative labels of the true null hypotheses are no longer uniform, i.e., $\mathbb{P}(L_{i}=1)=c$ and $\mathbb{P}(L_{i}=-1)=1-c$. In the proteomics community, the application of Equation (2) in the peptide detection problem is referred to as target-decoy competition or TDC for short. As such, we will refer to the method as TDC in the peptide detection setting, and SSS+ in the variable selection setting and more generally.
Alternatively, we can control the FDP by using, for example, FDP-stepdown (FDP-SD, Algorithm S2[40]. FDP-SD works by first reranking the labels $L$ according to the scores $W$ in descending order. Next, it compares the number of decoy (knockoff) wins in the first $i$ hypotheses, denoted as $D_{i}:=\#\{L_{j}=-1:j\leq i\}$, with a precomputed bound $\delta_{i}$, which is determined by $\alpha$ and $\gamma$. Each comparison is made in a ‘stepdown’ fashion, i.e., if the first $i$ hypotheses satisfy $D_{1}\leq\delta_{1},\dots,D_{i}\leq\delta_{i}$, then it moves on to the next comparison at $i+1$. For small $i$, it is impossible for $D_{i}\leq\delta_{i}$ and so FDP-SD begins this analysis at $i_{0}$, the smallest possible index for which $D_{i}\leq\delta_{i}$. Under Assumption 1, reporting the following list of discoveries $R_{\alpha,\gamma}$ controls the FDP:
$R_{\alpha,\gamma}:=\{i\in[p]:i\leq k_{FDP},L_{i}=1\},\quad k_{FDP}:=\max\{i:\Pi_{j=i_{0}}^{i}1_{D_{j}\leq\delta_{j}}=1\text{ or }i=0\}.$ (3)
@@ 2.1 Extension to side information
In proteomics, so-called post-processors such as PeptideProphet [28] and Percolator [26] have popularized the use of side information to increase the number of detected peptides. These tools use machine learning procedures, leveraging features associated with the PSMs, such as the charge state associated with the spectrum and the matched peptide’s length, to better distinguish correct from incorrect PSMs, and hence to deliver more discoveries. These tools have no theoretical guarantees and, having recently pointed out scenarios where such post-processors can fail to control the FDR [18], we will not analyze these tools in this paper.
To our knowledge, the only ‘intrinsic’ competition-based multiple testing procedure that uses arbitrary side information in a data-driven manner is the recent Adaptive Knockoffs (AdaKO) by Ren and Candès [44]. AdaKO is a specialization of the p-value based procedure AdaPT [33] that we discuss in the next section. By ‘intrinsic’, we mean that while other p-value based approaches may be modified to be used in the competition setting, they are significantly less powerful and do not really qualify in this sense, as demonstrated in [44]. Adaptive Knockoffs works by iteratively pruning $\mathcal{S}$, the candidate set of hypotheses, selecting each time the hypothesis it deems most likely to be a true null. To do that, it fits a classification algorithm using a strict set of information that partially masks the data: the scores $W$, the side information $\mathbf{x}$, and the labels $L_{i}$ of the hypotheses that were pruned up to that point ($i\in[p]\setminus\mathcal{S}$), as well as #$\#\{L_{i}=1:i\in S\}$ and #$\#\{L_{i}=-1:i\in S\}$. In their paper, Ren and Candès investigate this pruning strategy using different types of ‘filters’: a logistic regression (LR), a generalized additive model (GAM) which is the default filter in their R package, a random forest (RF), and a two-group probabilistic model which is fitted using the expectation-maximization algorithm (EM) [13]. The pruning process terminates when the estimated FDR is less than or equal to $\alpha$:
$\widehat{FDR}_{\mathcal{S}}:=\frac{\#\{i\in\mathcal{S},L_{i}=-1\}+1}{\#\{i\in\mathcal{S},L_{i}=1\}\vee 1}\leq\alpha.$ (4)
The remaining positively labeled hypotheses in $\mathcal{S}$ are then reported. It was shown that, given the following natural extension of Assumption 1, Adaptive Knockoffs controls the FDR.
@@
Let $N$ be the indices of the true null hypotheses. The labels $\{L_{i}:i\in N\}$ are i.i.d. ±$\pm 1$ uniform random variables independent of all the scores $W$, the side information $\mathbf{x}$, and the labels of the false null hypotheses.
@@ 3 Background: p-value based multiple testing
Multiple testing with FDR control was popularized by Benjamini and Hochberg when they developed the first testing procedure of this kind, which we refer to as the ‘BH procedure’ [2]. Since then, many p-value based methods that control the FDR have been developed to improve statistical power.
One such strategy for improving on BH involves estimating the fraction of true null hypotheses, denoted as $\pi_{0}$. Indeed, the BH procedure controls the FDR conservatively at level $\pi_{0}\cdot\alpha$ instead of $\alpha$. Hence, in cases when $\pi_{0}\ll 1$, the BH procedure loses a considerable amount of power. Methods like those by Storey et al. [46] and Benjamini et al. [3] estimate $\pi_{0}$, and essentially ‘plug-in’ the estimate $\hat{\pi}_{0}$, by applying the BH procedure at a threshold of $\alpha/\hat{\pi}_{0}$ instead of $\alpha$.
Instead of the FDR, Guo and Romano’s stepdown approach [20], which we refer to as GR-SD, controls the FDP. In fact, this stepdown approach inspired the development of FDP-SD in the competition setting. GR-SD works by first sorting the p-values in ascending order, i.e., $p_{1}\leq p_{2}\leq...\leq p_{m}$. Then it constructs a sequence of thresholds, $\delta_{i}$ for each p-value $p_{i}$ and searches for the largest $k$ such that $p_{1}\leq\delta_{1},\dots,p_{k}\leq\delta_{k}$ mutually holds, and then reports the corresponding $k$ smallest p-values. We provide the pseudocode for this method in Algorithm S3.

### v3

@@ Abstract
Standard multiple testing procedures are designed to report a list of discoveries, or suspected false null hypotheses, given the hypotheses’ p-values or test scores. Recently there has been a growing interest in enhancing such procedures by combining additional information with the primary p-value or score. In line with this idea, we develop RESET (REScoring via Estimating and Training), which uses a unique data-splitting protocol that subsequently allows any semi-supervised learning approach to factor in the available side information while maintaining finite sample error rate control. Our practical implementation, RESET Ensemble, selects from an ensemble of classification algorithms so that it is compatible with a range of multiple testing scenarios without the need for the user to select the appropriate one. We apply RESET to both p-value and competition based multiple testing problems and show that RESET is (1) power-wise competitive, (2) fast compared to most tools and (3) able to uniquely achieve finite sample false discovery rate or false discovery exceedance control, depending on the user’s preference.
Corresponding authors: jack.freestone@mq.edu.au, uri.keich@sydney.edu.au
@@ 1 Introduction
Scientists are often interested in testing many null hypotheses, $\{H_{i}:i=1,\dots,m\}$. Large-scale hypothesis testing is usually achieved by assigning a p-value, $p_{i}$, or a score, $Z_{i}$, to each null hypothesis. The collection of p-values, or scores, then undergoes a filtering process that rejects a subset of these null hypotheses subject to type-1 error rate control to minimize the number of misreported true null hypotheses.
@@
The most common choice of error rate is the false discovery rate (FDR), which is defined as the expectation of the false discovery proportion (FDP) — the fraction of true null hypotheses in the reported list of rejections, or discoveries (Benjamini and Hochberg (1995)). In this case, the filtering procedure ensures that the FDR is $\leq\alpha$ where $\alpha\in\left(0,1\right)$ is a prespecified threshold. Alternatively, we might be interested in controlling false discovery exceedance (FDX), which is the upper tail probability of the FDP. The FDX is said to be controlled with threshold $\alpha$ at confidence $1-\gamma$ if $\mathbb{P}(FDP>\alpha)\leq\gamma$. Evidently, such FDX control reduces the chance that the realized FDP exceeds $\alpha$, which is not guaranteed by FDR control (Guo and Romano (2007); Katsevich and Ramdas (2020); Goeman et al. (2021); Li et al. (2022); Luo et al. (2023); Ebadi et al. (2023)).
@@
Recently, researchers have focused on developing methods that leverage side information $\mathbf{x}_{i}\in\mathcal{X}$, which complements each p-value, or score, to better discriminate between true and false null hypotheses. These methods utilize a variety of strategies but ultimately operate with the same goal in mind: the side information is used to weight or rescore the hypotheses according to how confidently we believe that each hypothesis is a false null.
@@
In this paper, we offer a novel approach for multiple testing with side information. The idea is to randomly split a set of suspected true nulls into two sets, where one set is used for training a semi-supervised learning model to discriminate between the true and false null hypotheses, and the other for estimating the number of false discoveries. Accordingly, we refer to our new approach as RESET—REScoring via Estimating and Training. Our method flexibly incorporates any semi-supervised approach to help distinguish between true and false null hypotheses. Moreover, it is applicable to both competition- and p-value-based multiple testing scenarios, and allows the user to choose between FDR or FDX control. We evaluate RESET in a range of simulations and real data applications to verify that it is a competitive alternative to existing methods in terms of statistical power and speed.
@@ 2 Background: competition-based multiple testing
In competition-based multiple testing, each hypothesis $H_{i}$ is associated with two scores: a test or target score, $Z_{i}$, where larger values indicate greater evidence against the null, and a decoy score, $\tilde{Z}_{i}$, a random draw from the null distribution. Each pair $(Z_{i},\tilde{Z}_{i})$ is evaluated in a head-to-head competition yielding a label $L_{i}\in\{\pm 1\}$ that indicates which score is higher and a winning score $W_{i}$:
$L_{i}:=\begin{cases}+1&Z_{i}>\tilde{Z}_{i}\\ -1&Z_{i}<\tilde{Z}_{i}\end{cases},\quad W_{i}:=f(Z_{i},\tilde{Z}_{i}),$
where $f$ is symmetric in its arguments and increases with the values of $Z_{i}$ and $\tilde{Z}_{i}$, or with the difference between them, e.g., $f(x,y)=\max\{x,y\}$ or $f(x,y)=\lvert x-y\rvert$. Hypotheses with ties (i.e., $Z_{i}=\tilde{Z}_{i}$) are either discarded or the ties are randomly broken. We refer to a hypothesis with $L_{i}=1$ as a target win or a target for short, and $L_{i}=-1$ as a decoy win or a decoy for short. The winning scores and labels have two properties. First, the false nulls with large $Z_{i}$ will likely win their competition, resulting in a target win ($L_{i}=1$) with large $W_{i}$. Second, the true null labels are independent and equally likely to be $\pm 1$, independently of everything else. This is formalized in the following assumption.
@@
Let $N$ be the indices of the true null hypotheses. Conditioned on all the scores $W$, and the labels of the false null hypotheses, the true null labels $\{L_{i}:i\in N\}$ are i.i.d. $\pm 1$ uniform random variables.
Given Assumption 1 holds, the number of decoy wins ($L_{i}=-1$) that score above a cutoff $\tau$ can be used to conservatively estimate the number of true null hypotheses with $L_{i}=1$ that are also above $\tau$. Using this strategy, we can report a list of discoveries in one of two ways.
The first and by far the more common way is to apply Selective SeqStep+ (SSS+, Algorithm A1) (Barber and Candès (2015)), which controls the FDR under Assumption 1 at a prespecified threshold $\alpha$. Assuming the labels $L$ are reranked according to the scores $W$ in descending order, SSS+ reports the following list of discoveries, $R_{\alpha}$:
$R_{\alpha}:=\{i\leq\tau:L_{i}=1\},\text{ where }\tau:=\max\left\{t\in[m]:\frac{\#\{i\leq t,L_{i}=-1\}+1}{\#\{i\leq t,L_{i}=1\}\vee 1}\leq\alpha\right\},$ (1)
where $m]:=\{1,2,\dots,m\}$. Notably, SSS+ can be adjusted to handle a more general setting when the probabilities of the positive and negative labels of the true null hypotheses are no longer uniform (Barber and Candès (2015); Lei and Fithian (2016)).
Alternatively, we can control the FDX by using, for example, FDP-stepdown (FDP-SD, Algorithm A2) (Luo et al. (2023)). FDP-SD works by first reranking the labels $L$ according to the scores $W$ in descending order. Next, it compares the number of decoy wins in the first $i$ hypotheses, denoted as $D_{i}:=\#\{L_{j}=-1:j\leq i\}$, with a precomputed bound $\delta_{i}$, which is determined by $\alpha$ and $\gamma$. The comparisons are made in a ‘stepdown’ fashion: we test whether $D_{i}\leq\delta_{i}$ only if $D_{j}\leq\delta_{j}$ for $j<i$. For small $i$, it is impossible for $D_{i}\leq\delta_{i}$ and so FDP-SD begins this analysis at $i_{0}$, the smallest possible index for which $D_{i}\leq\delta_{i}$ can occur. Under Assumption 1, reporting the following list of discoveries $R_{\alpha,\gamma}$ controls the FDX:
$R_{\alpha,\gamma}:=\{i\in[m]:i\leq k_{FDP},L_{i}=1\},\quad k_{FDP}:=\max\{i:\Pi_{j=i_{0}}^{i}1_{D_{j}\leq\delta_{j}}=1\text{ or }i=0\}.$ (2)
@@ 2.1 Extension to side information
To our knowledge, the only ‘intrinsic’ competition-based multiple testing procedure that uses arbitrary side information in a data-driven manner is the recent Adaptive Knockoffs (AdaKO) by Ren and Candès (2023). AdaKO is a specialization of the p-value based procedure AdaPT (Lei and Fithian (2018)) that we introduce in the next section with further details in Section C.1. By ‘intrinsic’, we mean that while other p-value based approaches may be modified to be used in the competition setting, Ren and Candès (2023) demonstrate they are substantially less powerful, and are not viable in this sense. Adaptive Knockoffs works by iteratively pruning $\mathcal{\hat{S}}_{t}$, the candidate set of hypotheses at each step $t$, by selecting the hypothesis it deems most likely to be a true null. To do that, it fits a classification algorithm using a strict subset of information that partially masks the data: the scores $W$, the side information $\mathbf{x}$, and the labels $(L_{i})_{i\in[m]\setminus\mathcal{\hat{S}}_{t}}$ of the hypotheses that were pruned up to that point, as well as $\{i\in\mathcal{\hat{S}}_{t}:L_{i}=1\}$ and $\{i\in\mathcal{\hat{S}}_{t}:L_{i}=-1\}$. In their paper, Ren and Candès (2023) investigate this pruning strategy using different types of filters: a logistic regression (LR), a generalized additive model (GAM) which is the default filter in their R package, a random forest (RF), and a two-group model which is fitted using the expectation-maximization algorithm (EM) (Dempster et al. (1977)). The pruning process terminates at step $\tau$ when the estimated FDR is less than or equal to $\alpha$:
$\widehat{FDR}_{\mathcal{\hat{S}}_{\tau}}:=\frac{\#\{i\in\mathcal{\hat{S}}_{\tau}:L_{i}=-1\}+1}{\#\{i\in\mathcal{\hat{S}}_{\tau}:L_{i}=1\}\vee 1}\leq\alpha.$ (3)
The remaining positively labelled hypotheses in $\mathcal{\hat{S}}_{\tau}$ are then reported. It was shown that, given the following natural extension of Assumption 1, Adaptive Knockoffs controls the FDR.
@@
Let $N$ be the indices of the true null hypotheses. Conditioned on all the scores $W$, the side information $\mathbf{x}$, and the labels of the false null hypotheses, the true null labels $\{L_{i}:i\in N\}$ are i.i.d. $\pm 1$ uniform random variables.
@@ 3 Background: p-value based multiple testing
Multiple testing with FDR control was popularized by the BH procedure (Benjamini and Hochberg (1995)). Since then, numerous p-value-based methods have been developed to address multiple testing in two main ways. First, many methods build on the BH framework to improve statistical power by estimating the fraction of true null hypotheses denoted as $\pi_{0}$ (Storey et al. (2004); Benjamini et al. (2006); Hu et al. (2010)), reweighting p-values (Genovese et al. (2006); Blanchard and Roquain (2008); Roquain and Van De Wiel (2009); Habiger (2017); Ramdas et al. (2019)), or by reordering them (Barber and Candès (2015); Lei and Fithian (2016); G’Sell et al. (2016); Li and Barber (2017)). A second approach focuses on the stricter criterion of FDX control. One example of this second direction is the stepdown procedure proposed by Guo and Romano (2007), which we will refer to as GR-SD in our later analyses.